Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
Sposito - 14:56 Sunday 17 December 2023 (62766) Print this report
Operator Report - Morning shift

Today, I found the ITF in LOCKING_ARMS_IR with the north arm misaligned. I realigned it, moving the NE mirror. I restored the lock and performed a CITF alignment, and I was able to reach LOW_NOISE_3 on the first attempt.

After 30 minutes, as requested by I. Nardecchia,  07:40 UTC I set the new values for the NE and WE ring heaters with a common step of NE: -0.26 and WE: -0.21

  • 2023-12-17 07h40m22 UTC    sposito 'Tower1:TransientOFF' sent to RingHeaterNE
  • 2023-12-17 07h40m28 UTC    sposito 'Tower1:TransientOFF' sent to RingHeaterWE
  • 2023-12-17 07h40m56 UTC    sposito 'Tower1:PowerFactor' sent to RingHeaterNE
  • 2023-12-17 07h41m17 UTC    sposito 'Tower1:PowerFactor' sent to RingHeaterWE
  • 2023-12-17 07h41m25 UTC    sposito 'Tower1:TransientON' sent to RingHeaterNE
  • 2023-12-17 07h41m28 UTC    sposito 'Tower1:TransientON' sent to RingHeaterWE

13:50 UTC as agreeded with I. Nardecchia, I set the new values for the NE and WE ring heaters with a common step of NE: -0.15 and WE: +0.11

  • 2023-12-17 13h50m54 UTC    sposito 'Tower1:TransientOFF' sent to RingHeaterNE
  • 2023-12-17 13h50m58 UTC    sposito 'Tower1:TransientOFF' sent to RingHeaterWE
  • 2023-12-17 13h51m21 UTC    sposito 'Tower1:PowerFactor' sent to RingHeaterNE
  • 2023-12-17 13h51m29 UTC    sposito 'Tower1:PowerFactor' sent to RingHeaterWE
  • 2023-12-17 13h51m36 UTC    sposito 'Tower1:TransientON' sent to RingHeaterNE
  • 2023-12-17 13h51m40 UTC    sposito 'Tower1:TransientON' sent to RingHeaterWE

ITF left in LOW_NOISE_3

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×